OpenStudy (anonymous):

\[\int\limits_{1}^{3} (x/(x^2+3))\]

OpenStudy (anonymous):

use substiution \[u =x^2+3\] then \[du =2x dx\] =>\[(1/2)du=xdx\] changing bounds upper \[u=3^3+3=12\]lower, \[u=1^2+3=4\] now we have \[\int\limits_{4}^{12} ((1/2)u)du\] \[=(1/2)\int\limits\limits_{4}^{12} (u)du\] \[=(1/2)[ u^2/2]|_{4}^{12}\] \[=(1/2)[ 12^2/2 - 4^2/2]\] \[=32\]

OpenStudy (anonymous):

\[\int\limits_{1}^3{\frac{x}{x^2+3}}dx\] \[\text{Let }u=x^2+3\].....

OpenStudy (anonymous):

are you sure? when I do it on my calculator i get the answer (ln(3))/2 as a answer. it is the same answer i get from using wolframalpha too.

OpenStudy (anonymous):

my problem is that i have know idea how to set it up manually

OpenStudy (anonymous):

you might be right 1 min

OpenStudy (anonymous):

sorry I set up the intergal worng it should be this \[\int\limits_{4}^{12} (1/2)(1/u)du\] \[ =(1/2)\int\limits\limits_{4}^{12}(1/u)du\] \[ =(1/2)[\ln u]|_{4}^{12}\] \[ =(1/2)[\ln 12 -\ln4]\] \[ =(\ln 3)/2\]

OpenStudy (anonymous):

no problem. Thanks :)